# Positioning-Aided Channel Estimation for Multi-LEO Satellite Downlink Communications

### v1

@@ Abstract
We investigate a multi-low Earth orbit (LEO) satellite system that simultaneously provides positioning and communication services to terrestrial user terminals. To address the challenges of channel estimation in LEO satellite systems, we propose a novel two-timescale positioning-aided channel estimation framework, exploiting the distinct variation rates of position-related parameters and channel gains inherent in LEO satellite channels. Using the misspecified Cramér-Rao bound (MCRB) theory, we systematically analyze positioning performance under practical imperfections, such as inter-satellite clock bias and carrier frequency offset. Furthermore, we theoretically demonstrate how position information derived from downlink positioning can enhance uplink channel estimation accuracy, even in the presence of positioning errors, through an MCRB-based analysis. To overcome the constraints of limited link budgets and communication rates associated with single-satellite-based communication, we develop a distributed beamforming strategy for downlink communication. This strategy allows LEO satellites to independently optimize their beamformers using local channel state information, eliminating the need for centralized processing while preserving the advantages of multi-satellite cooperative communication. Theoretical analyses and numerical results confirm the effectiveness of the proposed framework in achieving high-precision downlink positioning under practical imperfections, facilitating uplink channel estimation, and enabling efficient downlink communication.
@@
LEO satellite positioning and communication, channel estimation, MCRB, distributed beamforming.
@@ I Introduction
The evolution of mobile communications has been driven by a remarkable vision: enabling seamless, high-speed information access from any point on Earth. While generational advances in cellular networks, from 1G to 5G and beyond, have largely achieved this goal in urban and suburban areas where infrastructure deployment is economically viable, significant challenges remain in rural and remote regions. In these underdeveloped areas, reliable communication and internet access often remain either unavailable or prohibitively expensive, contributing to a growing digital divide[1]. The non-terrestrial networks (NTN)-based communications have emerged as a promising solution to bridge this gap, with airborne and spaceborne platforms such as high-altitude platform stations and satellites, serving as access points to enable truly global connectivity[2]. Among these candidates, low Earth orbit (LEO) satellites have attracted particular attention due to their distinct advantages over medium Earth orbit (MEO) and geostationary Earth orbit (GEO) satellites, including reduced path loss and propagation delay, greater constellation flexibility, and lower launch costs[3, 4, 5].
@@
Although satellite channels pose significant challenges for CSI acquisition, they exhibit distinct characteristics compared to terrestrial channels, particularly stronger line-of-sight (LOS) dominance and geometric properties [9]. These characteristics are inherently related to position information [5], suggesting that strategic utilization of UT position information could potentially address channel estimation challenges in LEO satellite communications, especially by enabling partial channel inference through the reconstruction of components such as steering vectors. However, this approach raises critical research questions, such as how to achieve accurate and efficient position acquisition within LEO satellite communication networks and how to effectively leverage this position information to enhance LEO satellite channel estimation. On the other hand, to address the limited link budget of single-satellite systems, multi-satellite cooperative transmission has emerged as a promising solution [10, 11, 12, 13, 14, 15]. However, achieving coherent transmission across satellites within non-orthogonal frequency or time resources presents significant challenges, such as phase synchronization among different satellites and the need for centralized processing based on global CSI[10]. These centralized operations introduce substantial processing overhead and latency, further complicating the practical implementation of cooperative transmission schemes.
@@
Concurrent with advances in LEO satellite communication, there is growing interest in exploring LEO satellites as potential supplements or alternatives to GNSS [22, 23]. This interest arises from the inherent advantages of LEO satellites over GNSS, including stronger signal reception, extensive constellations, and broad frequency diversity, characteristics that have also partially driven the success of LEO satellite communications [23, 24, 25]. Recent research has begun to investigate the integration of LEO satellite communication and positioning functionalities. For instance, [26] proposes a massive multi-input-multi-output (MIMO) LEO satellite system that supports simultaneous communication and positioning in the downlink, designing hybrid beamforming to meet their distinct requirements. However, this work primarily focuses on the tradeoff between positioning and communication performance, overlooking their mutual benefits such as the potential of positioning in facilitating channel estimation, which is a critical challenge in LEO satellite systems. Moreover, it does not account for the inevitable clock bias and carrier frequency offset (CFO) among LEO satellites, factors that may significantly affect positioning performance.
Three key questions emerge under this context: 1) how can the positioning performance of LEO satellites be effectively characterized under practical imperfections such as clock bias and CFO; 2) how can the position information of UTs obtained through LEO satellite positioning be leveraged to assist the challenging channel estimation task; and 3) how can the benefits of the improved link budget offered by multi-satellite cooperative communication be harnessed while avoiding the complexities of synchronization and centralized processing?
In this paper, we investigate a scenario where a cluster of LEO satellites, each equipped with a uniform planar array (UPA), simultaneously provides downlink positioning and communication services to multiple single-antenna terrestrial UTs. We aim to: 1) characterize the positioning performance under practical impairments such as clock bias/CFO incurred by imperfect time-/frequency-domain synchronization; 2) unveil how the UTs’ position information obtained by downlink positioning can be exploited to enhance uplink channel estimation; and 3) explore distributed beamforming relying solely on local CSI acquired at each satellite, enabling efficient downlink communication while avoiding the overhead and complexity typically associated with centralized operations. The main contributions of this paper are as follows.
@@
  • Building on the distinct variation rates of position-related parameters and random channel gains in LEO satellite channels, as detailed in Section II, we propose a two-timescale frame structure. Downlink positioning operates on the longer position-coherent timescale, while uplink channel estimation and downlink communication occur on the shorter channel gain-coherent timescale. Through this framework, each process is scheduled according to its variation rate, maximizing resource utilization while preserving the potential to leverage positioning for efficient channel estimation and communication.
  • To characterize the LEO satellite-based positioning performance under imperfect inter-satellite clock bias and CFO, we employ misspecified Cramér-Rao bound (MCRB) theory rather than traditional Cramér-Rao bound (CRB) analysis. The MCRB framework is specifically designed to derive performance limits for estimation under model mismatch[27]. Through this theoretical approach, we rigorously analyze how imperfect inter-satellite clock bias and CFO affect time-of-arrival (TOA) and Doppler shift measurements, and systematically derive their impact on positioning accuracy.
  • We demonstrate the advantage of leveraging the position information of UTs, obtained during the positioning stage, to enhance uplink channel estimation. Specifically, we account for the inevitable positioning errors, which introduce model mismatches in the positioning-aided channel estimation (PACE) process. By invoking the MCRB theory, we provide a theoretical characterization of the uplink channel estimation performance facilitated by the UTs’ position information.
  • We propose a distributed beamforming strategy for downlink communication, where each UT is served by the LEO satellites withn the cluster simultaneously. In this approach, each LEO satellite independently optimizes its beamformer using only its local CSI with respect to UTs, thereby eliminating the need for a centralized processor and complex inter-satellite coordination mechanisms.
@@
This paper is structured as follows. We begin by presenting the system model in Section II. Section III examines both downlink positioning and uplink channel estimation performance through MCRB theory, incorporating the PACE. A distributed beamforming approach for downlink communications is detailed in Section IV. Numerical results are presented in Section V, with conclusions drawn in Section VII.
@@ II System Model
As illustrated in Fig. 1, we consider a scenario where a cluster of $S$ LEO satellites serves $U$ UTs, simultaneously providing downlink positioning and communication services. Each LEO satellite is equipped with a UPA comprising $N=N_{\text{h}}\times N_{\text{v}}$ half-wavelength-spaced antennas, where $N_{\text{h}}$ and $N_{\text{v}}$ denote the numbers of antennas in the horizontal and vertical dimensions, respectively. Each UT is equipped with a single antenna. While accurate CSI is essential for achieving efficient downlink communication, channel estimation is challenging in space-borne communication systems due to the limited link budget and short coherence time. These challenges originate from the long travel distance between the satellites and the UTs as well as the high-speed satellite movement[7]. In the following, we first introduce the adopted channel model. Subsequently, we elaborate on a two-timescale property of this model, which can be leveraged to facilitate efficient channel acquisition at LEO satellites, thereby enhancing downlink communications.
@@ II-A Channel Model
We consider a satellite communication system operating at relatively high-frequency bands such as the Ku and Ka bands, where the channel is dominated by LOS propagation[9, 13]. [^1] While the multipath effect can still play a significant role in dense urban environments, LEO satellite communications are primarily designed to serve suburban, rural, and remote areas, such as oceans and deserts, where terrestrial infrastructure is sparse or nonexistent, and scatterers are greatly reduced. Consequently, LOS propagation becomes even more dominant. Assume the system employs the orthogonal frequency division multiplexing (OFDM) scheme with $K$ subcarriers. Let $\Delta f=B/K$ and $T=1/B$ denote the subcarrier spacing and symbol duration, respectively, where $B$ is the signal bandwidth. The channel from the $s$-th LEO satellite to the $u$-th UT during the $\ell$-th symbol over the $k$-th subcarrier is expressed as [^2] Note that, though not further indexed for symbolic simplicity, this expression is viewed within one channel-gain coherent interval (as elaborated later) and $\upsilon_{s,u}$, $\tau_{s,u}$, and $\bm{\theta}_{s,u}$ are viewed constant for many such intervals.
@@
where $\alpha_{s,u}$ is the complex channel gain, $G(\cdot)$ is the antenna radiation gain determined solely by the elevation angle[28], $\tau_{s,u}$ and $\upsilon_{s,u}$ represent the TOA and Doppler shift, respectively, and $\mathbf{a}(\bm{\theta}_{s,u})\in\mathbb{C}^{N}$ is the steering vector at the LEO satellite, with $\bm{\theta}_{s,u}=[\theta_{s,u}^{\mathrm{az}},\theta_{s,u}^{\mathrm{el}}]^{\mathsf{T}}$ being the angle-of-departure (AOD) composed of both azimuth and elevation angles. Without loss of generality, we assume the UPA at each satellite is deployed on the XY-plane of its local coordinate system, as illustrated in Fig. 2. Hence, the steering vector can be expressed as
@@
[…]
@@
Figure 2: Geometric configuration of the local coordinate systems for LEO satellites, depicting the UPA deployment in the XY-plane.
@@ II-B1 Timescale Analysis
  • Large-Timescale Parameters: These parameters are associated with the UT’s position, including the Doppler shift, TOA, and AOD. Due to the large geographical scale of satellite communications, they typically change slowly (on timescales of seconds) or can be accurately predicted through tracking mechanisms to reduce estimation overhead[25]. Although LEO satellites move rapidly, their orbits are typically predictable and well-determined. Therefore, while these parameters are related to the fast-varying position of LEO satellites, they can still be considered slow-varying due to their predictability.
  • Small-Timescale Parameters: These parameters consist of random variations in the amplitude and phase of the channel gain. In contrast to the slow-varying parameters determined by geometry, these parameters are inherently unpredictable and fluctuate rapidly. For instance, tropospheric scintillation can cause phase and amplitude variations in Ka-band LEO satellite channels that fluctuate on sub-second timescales, with the exact rate depending on atmospheric and orbital conditions[31].
Notably, the Doppler shift caused by the high-speed movement of the satellite induces rapid temporal phase rotation in the channel. However, this effect can be largely mitigated through carrier synchronization at the receiver, albeit with minor residual errors [3]. As a result, its impact on LEO satellite communications is limited. Nonetheless, we retain it in the channel model because it can serve as a valuable measurement for positioning, as will be discussed later. For positioning purposes, the Doppler shift should be explicitly estimated before being compensated for communications.
@@ II-B2 Two-Timescale Frame Structure
Inspired by the timescale analysis, we propose a two-timescale frame structure that exploits the inherent two-timescale property of LEO channels to enable efficient channel estimation/reconstruction, thus facilitating downlink LEO transmission. As shown in Fig. 3, the duration of a large-timescale frame is determined by the position-coherent time, which indicates the interval after which the UTs’ position information becomes outdated. Each large-timescale frame begins with an initial subframe for downlink positioning (where position estimates obtained by UTs are fed back to LEO satellites), followed by multiple small-timescale subframes during which the channel gain remains coherent. Each small-timescale subframe is further divided into slots for uplink channel estimation and downlink communication. Note that time division duplex (TDD) is adopted to differentiate between uplink and downlink transmissions. The proposed approach leverages uplink-downlink channel reciprocity, which applies uplink channel estimation to optimize downlink communication design.
@@
In the following, we will first conduct a performance analysis based on the two-timescale PACE and then propose a distributed beamforming method across multiple LEO satellites for downlink communication.

### v4

# Positioning-Aided Channel Estimation for Multi-LEO Satellite Cooperative Beamforming

@@ Abstract
We investigate a multi-low Earth orbit (LEO) satellite system that simultaneously provides positioning and communication services to terrestrial user terminals. To address the challenges of accurately acquiring channel state information in LEO satellite systems, we propose a novel two-timescale positioning-aided channel estimation framework, exploiting the distinct variation rates of position-related parameters and channel gains inherent in LEO satellite channels. Using the misspecified Cramér-Rao bound (MCRB) theory, we systematically analyze positioning performance under practical imperfections, such as inter-satellite clock bias and carrier frequency offset. Furthermore, we theoretically demonstrate how position information derived from downlink positioning can enhance uplink channel estimation accuracy, even in the presence of positioning errors, through an MCRB-based analysis. To address the limited link budgets and communication rates of single-satellite communication, we develop a multi-LEO cooperative beamforming strategy for downlink transmission that leverages cluster-wise satellite cooperation while maintaining reduced complexity. Theoretical analyses and numerical results confirm the effectiveness of the proposed framework in facilitating high-precision downlink positioning under practical imperfections, facilitating uplink channel estimation, and enabling efficient downlink communication.
@@
LEO satellite positioning and communication, channel estimation, MCRB, cooperative beamforming.
@@ I Introduction
The evolution of mobile communications has been driven by a remarkable vision: enabling seamless, high-speed information access from any point on Earth. While generational advances in cellular networks, from 1G to 5G and beyond, have largely achieved this goal in urban and suburban areas where infrastructure deployment is economically viable, significant challenges remain in rural and remote regions. In these underdeveloped areas, reliable communication and internet access often remain either unavailable or prohibitively expensive, contributing to a growing digital divide[1]. The non-terrestrial networks (NTN)-based communications have emerged as a promising solution to bridge global connectivity gaps, leveraging airborne and spaceborne platforms, such as high-altitude platforms (HAPs) and satellites, as aerial access points[2, 3, 4, 5, 6, 7]. Among these platforms, low Earth orbit (LEO) satellites have garnered particular attention due to several advantages over medium Earth orbit (MEO), geostationary Earth orbit (GEO) satellites, and HAPs. These advantages include lower path loss and reduced propagation delays, increased flexibility in constellation design, and comparatively lower deployment and launch costs[8, 9, 10].
@@
Although satellite channels pose significant challenges for CSI acquisition, they exhibit distinct characteristics compared to terrestrial channels, particularly strong line-of-sight (LOS) dominance and pronounced geometric structures [14]. These properties are intrinsically tied to position information [10], suggesting that judicious use of UT position knowledge could help alleviate channel estimation challenges in LEO satellite communications, e.g., by enabling partial channel inference through the reconstruction of geometry-dependent components such as steering vectors. This viewpoint naturally raises two key questions: (i) how to acquire accurate UT position information efficiently in LEO satellite networks, and (ii) how to leverage such information to improve LEO channel estimation and, ultimately, communication performance. On the other hand, to overcome the limited link budget inherent in single-satellite service, multi-satellite cooperative transmission has emerged as a promising solution under diverse design objectives [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Recent works have, for example, designed cooperative transmission schemes to proactively mitigate asynchronous interference in single-carrier/statistical-CSI settings [22], extended such treatments to orthogonal frequency division multiplexing (OFDM) systems (often resorting to high-complexity quadratically constrained quadratic program (QCQP)-based formulations) [23], or emphasized joint power allocation and user scheduling with heuristic precoder constructions [24]. While these advances are highly relevant, they are not directly aligned with the objective of this paper, which uses cooperative beamforming primarily as a controlled baseline to assess the utility of positioning-aided CSI. Moreover, despite the growing interest in multi-satellite cooperation, the computational scalability of cooperative beamforming remains a key bottleneck, especially for resource-constrained LEO satellite platforms.
@@
Concurrent with advances in LEO satellite communication, there is growing interest in exploring LEO satellites as potential supplements or alternatives to GNSS [30, 31, 32, 33, 34]. This interest arises from the inherent advantages of LEO satellites over GNSS, including stronger signal reception, extensive constellations, and broad frequency diversity, characteristics that have also partially driven the success of LEO satellite communications [31, 35, 36]. Recent research has begun to investigate the integration of LEO satellite communication and positioning functionalities. For instance, [37] proposes a massive multi-input-multi-output (MIMO) LEO satellite system that supports simultaneous communication and positioning in the downlink, designing hybrid beamforming to meet their distinct requirements. However, this work primarily focuses on the tradeoff between positioning and communication performance, overlooking their mutual benefits such as the potential of positioning in facilitating channel estimation, which is a critical challenge in LEO satellite systems. Moreover, it does not account for the inevitable clock bias and carrier frequency offset (CFO) among LEO satellites, factors that may significantly affect positioning performance.
In this paper, we investigate a cluster of LEO satellites, each equipped with a uniform planar array (UPA), that simultaneously provide downlink positioning and communication services to multiple single-antenna terrestrial UTs. We aim to address the key challenges that specifically arise in this setting: characterizing positioning performance under practical synchronization imperfections such as inter-satellite clock bias and CFO; leveraging user position information to enhance channel estimation while accounting for inevitable positioning errors; and mitigating single-satellite link-budget limitations through a carefully designed cooperative multi-satellite beamforming scheme with reduced optimization complexity.
The main contributions of this paper are as follows.
@@
  • Building on the distinct variation rates of position-related parameters and random channel gains in LEO satellite channels, as detailed in Section II, we propose an innovative two-timescale frame structure. Downlink positioning operates on the longer position-coherent timescale, while uplink channel estimation and downlink communication occur on the shorter channel gain-coherent timescale. Through this framework, each process is scheduled according to its variation rate, maximizing resource utilization while preserving the potential to leverage positioning for efficient channel estimation and communication.
  • To move beyond idealized synchronization assumptions commonly adopted in prior works, we extend the use of the misspecified Cramér-Rao bound (MCRB) framework to explicitly capture the impact of inter-satellite clock bias and CFO. Unlike the classical Cramér-Rao bound (CRB), the MCRB is designed for estimation under model mismatch [38]. Through this lens, we rigorously quantify how these synchronization-related impairments distort time-of-arrival (TOA) and Doppler measurements and, in turn, degrade positioning accuracy in practical LEO scenarios.
  • Within the proposed positioning-aided channel estimation (PACE) framework, we show how large-timescale positioning errors propagate into small-timescale channel estimation, creating inherent model mismatch. By applying the MCRB, we systematically analyze how uncertainty in the positioning stage translates into performance loss in uplink channel estimation, thereby identifying the robustness margins required for practical position-aided designs.
  • We propose a cooperative beamforming baseline for downlink communication, where each UT is simultaneously served by all LEO satellites in a cluster. We adopt the weighted minimal mean squared error (WMMSE) framework and develop a low-complexity solution that strategically exploits block coordinate descent (BCD), strong duality, and matrix-inversion structures to mitigate the otherwise prohibitive computational burden of multi-satellite cooperation. Importantly, this baseline is designed to be computationally practical and to enable a transparent evaluation of the proposed PACE framework by comparing the performance achieved under nominal (PACE-estimated) versus ground-truth channel models.
@@
This paper is structured as follows. We begin by presenting the system model in Section II. Section III examines both downlink positioning and uplink channel estimation performance through MCRB theory, incorporating the PACE. A cooperative beamforming approach for downlink communications is detailed in Section IV. Numerical results are presented in Section V, with conclusions drawn in Section VI.
@@ II System Model
As illustrated in Fig. 1, we consider a scenario where a cluster of $S$ LEO satellites serves $U$ UTs, simultaneously providing downlink positioning and communication services. In this paper, we assume that no handovers occur during operation. The impact of both intra- and inter-cluster handovers is left for future investigation. Each LEO satellite is equipped with a UPA comprising $N=N_{\text{h}}\times N_{\text{v}}$ half-wavelength-spaced antennas, where $N_{\text{h}}$ and $N_{\text{v}}$ denote the numbers of antennas in the horizontal and vertical dimensions, respectively. Each UT is equipped with a single antenna. [^1] Note that extending single-antenna UTs to multi-antenna setting is a promising direction: receiver spatial resolution can help separate signals with distinct Dopplers and delays and mitigate asynchronous interference[23]. However, a rigorous treatment would also require modifying the positioning model to incorporate array orientation and angle-related measurements, which is beyond the present scope and will be pursued in future work. While accurate CSI is essential for achieving efficient downlink communication, channel estimation is challenging in space-borne communication systems due to the limited link budget and short coherence time. These challenges originate from the long travel distance between the satellites and the UTs as well as the high-speed satellite movement[12]. In the following, we first introduce the adopted channel model. Subsequently, we elaborate on a two-timescale property of this model, which can be leveraged to facilitate efficient channel acquisition at LEO satellites, thereby enhancing downlink communications. Note that handovers in LEO satellite systems typically occur on the order of tens of seconds to a few minutes, and thus span multiple position-coherent large-timescale frames (see further discussion later). Since the proposed PACE framework focuses on the integration of positioning and communication within each frame, we omit explicit handover modeling to avoid introducing an additional control-plane layer and associated redundancy beyond the scope of this work. This simplifying assumption is also widely adopted in prior studies such as[24, 19, 15, 16].
@@ II-A Channel Model
We consider a satellite communication system operating at relatively high-frequency bands such as the Ku and Ka bands, where the channel is dominated by LOS propagation[14, 18]. [^2] While the multipath effect can still play a significant role in dense urban environments, LEO satellite communications are primarily designed to serve suburban, rural, and remote areas, such as oceans and deserts, where terrestrial infrastructure is sparse or nonexistent, and scatterers are greatly reduced. Consequently, LOS propagation becomes even more dominant. Assume the system employs the OFDM scheme with $K$ subcarriers. Let $\Delta f=B/K$ and $T=1/\Delta f$ denote the subcarrier spacing and symbol duration, respectively, where $B$ is the signal bandwidth. The channel from the $s$-th LEO satellite to the $u$-th UT during the $\ell$-th OFDM symbol over the $k$-th subcarrier is expressed as [^3] Note that, though not further indexed for symbolic simplicity, this expression is viewed within one channel-gain coherent interval (as elaborated later) and $\upsilon_{s,u}$, $\tau_{s,u}$, and $\bm{\theta}_{s,u}$ are viewed constant for many such intervals.
@@
where $\alpha_{s,u}$ is the complex channel gain, $G(\cdot)$ is the symmetric antenna radiation pattern determined solely by the elevation angle[39], $\tau_{s,u}$ and $\upsilon_{s,u}$ represent the TOA and Doppler shift, respectively, and $\mathbf{a}(\bm{\theta}_{s,u})\in\mathbb{C}^{N}$ is the steering vector at the LEO satellite, with $\bm{\theta}_{s,u}=[\theta_{s,u}^{\mathrm{az}},\theta_{s,u}^{\mathrm{el}}]^{\mathsf{T}}$ being the angle-of-departure (AOD) composed of both azimuth and elevation angles. Without loss of generality, we assume the UPA at each satellite is deployed on the XY-plane of its local coordinate system. Hence, the steering vector can be expressed as
@@
$20\log_{10}\beta_{s,u}=\beta^{\text{FS}}_{s,u}+\beta^{\text{SF}}_{s,u}+\beta^{\text{CL}}_{s,u}+\beta^{\text{AB}}_{s,u}+\beta^{\text{SC}}_{s,u}\text{ [dB]},$ (4)
@@ II-B1 Timescale Analysis
  • Large-Timescale Parameters: These parameters are associated with the positions of the satellites and UTs, and include the Doppler shift, TOA, and AOD. Although LEO satellites move rapidly, their orbits are predictable and well-determined. Thus, the dominant deterministic Doppler component evolves coherently with other geometry-related parameters such as TOA and AOD. As long as the UTs’ positions do not change abruptly (which is typically the case in practice), this predictability remains valid over the position-coherent timescale, which is primarily governed by the update rate of the UTs’ positions rather than the satellites’ trajectories. Therefore, it is reasonable to assume that large-timescale parameters vary on the order of seconds or longer in typical scenarios.
  • Small-Timescale Parameters: These parameters consist of random variations in the amplitude and phase of the channel gain. Residual Doppler fluctuations that remain after compensation, due to estimation errors, are also absorbed into this category. Unlike the geometry-determined large-timescale parameters, these small-timescale effects are inherently unpredictable and fluctuate rapidly. Such variations can arise from tropospheric scintillation, which induces sub-second fluctuations in Ka-band LEO satellite channels depending on atmospheric and orbital conditions [42]. Moreover, even though the LOS component dominates at Ku/Ka bands, nearby mobile scatterers can still generate multipath components that merge with the LOS path, forming a composite channel gain that exhibits small-scale fading with amplitude and phase variations on millisecond timescales [28, 18].
Note that a similar dual-timescale effect has been analyzed and exploited in terrestrial systems [43, 27]. In LEO satellite channels, the Doppler shift induced by the high-speed motion of the satellite causes rapid temporal phase rotations. While this effect can be largely mitigated by carrier synchronization at the receiver, small residual errors may remain [8], limiting its impact on communication performance. Nevertheless, we retain the Doppler shift in the channel model since it also provides a valuable measurement for positioning, as will be discussed later. For positioning, the Doppler shift must be explicitly estimated prior to being compensated for communication purposes.
@@ II-B2 Two-Timescale Frame Structure
Inspired by the timescale analysis, we propose a two-timescale frame structure that exploits the inherent two-timescale property of LEO channels to enable efficient channel estimation/reconstruction, thus facilitating downlink LEO transmission. As shown in Fig. 2, the duration of a large-timescale frame is determined by the position-coherent time, which indicates the interval after which the UTs’ position information becomes outdated. Each large-timescale frame begins with an initial subframe for downlink positioning, where UTs’ position estimates can be fed back to LEO satellites [24, 44], followed by multiple small-timescale subframes during which the channel gain remains coherent. We would like to emphasize that the feedback carries only low-rate UT position information (a 3D real vector) that remains coherent over the large-timescale interval. This leads to negligible overhead and no stringent latency requirements. Moreover, the dominant uncertainty is the position estimation error (already modeled in our analysis), under which Fig. 5 shows that PACE-based channel reconstruction is robust in practical regimes. Therefore, explicit feedback-link errors are omitted. Note that time division duplex (TDD) [^4] While current NR-NTN bands (e.g., n255/n256 [45]) are mainly frequency division duplex (FDD), our PACE framework relies on uplink–downlink reciprocity and is therefore studied under a TDD assumption. Notably, TDD for LEO is explicitly considered by 3rd Generation Partnership Project (3GPP) [41], has operational precedent in systems such as Iridium Next, and is further examined in recent feasibility studies [46]. Hence, our TDD-based design complements the FDD baseline and provides forward-looking insights into reciprocity-enabled schemes for future NTN systems. is adopted to differentiate between uplink and downlink transmissions. The proposed approach leverages uplink-downlink channel reciprocity, which applies uplink channel estimation to optimize downlink communication design.
@@
In the following, we will first conduct a performance analysis based on the two-timescale PACE and then propose a cooperative beamforming method across multiple LEO satellites for downlink communication.